### v2

# Self-Labeling the Job Shop Scheduling Problem

Abstract
In this work, we propose a Self-Supervised training strategy specifically designed for combinatorial problems. One of the main obstacles in applying supervised paradigms to such problems is the requirement of expensive target solutions as ground-truth, often produced with costly exact solvers. Inspired by Semi- and Self-Supervised learning, we show that it is possible to easily train generative models by sampling multiple solutions and using the best one according to the problem objective as a pseudo-label. In this way, we iteratively improve the model generation capability by relying only on its self-supervision, completely removing the need for optimality information. We prove the effectiveness of this Self-Labeling strategy on the Job Shop Scheduling (JSP), a complex combinatorial problem that is receiving much attention from the Reinforcement Learning community. We propose a generative model based on the well-known Pointer Network and train it with our strategy. Experiments on popular benchmarks demonstrate the potential of this approach as the resulting models outperform constructive heuristics and current state-of-the-art learning proposals for the JSP.  [^1] The code is available at: https://github.com/AndreaCorsini1/SelfLabelingJobShop
1 Introduction
The Job Shop Problem is a classic optimization problem with many practical applications in industry and academia [51]. At its core, the JSP involves scheduling a set of jobs onto a set of machines, where each job has to be processed on the machines following a specific order. The goal of this problem is generally to minimize the makespan, i.e., the time required to complete all jobs [39].
Over the years, various approaches have been developed to tackle the JSP. A common one is adopting exact methods such as Mixed Integer Programming solvers (MIP). However, these methods often fail to provide either optimal or high-quality solutions on medium and large instances [29]. As a remedy, metaheuristics have been extensively investigated [35, 13, 24], and constitute an alternative to exact methods. Although state-of-the-art metaheuristics like [36] can provide quality solutions in a fair amount of time, typically minutes, they are rather complex to implement and is hard to reproduce their results [6]. Therefore, due to their lower complexity, and under tighter time constraints, Priority Dispatching Rules (PDR) [39] are often preferred, despite their lower quality. The main issue of PDRs is their tendency to perform well in some cases and poorly in others, primarily due to their blind prioritization schema based on rigid hardcoded and hand-crafted rules [21].
Therefore, recent works have started to investigate Machine Learning (ML) to enhance or replace some of these approaches, primarily focusing on PDRs. ML-based approaches, specifically Reinforcement Learning (RL) ones, demonstrated to deliver higher-quality solutions than classic PDRs at the cost of a small increase in execution time [32]. Despite the potential of RL [49, 25], training RL agents is a complex task [42, 33], is sensitive to hyper-parameters [40], and has reproducibility issues [22].
Contrary, supervised learning is less affected by these issues but heavily relies on expensive annotations [31]. This is particularly problematic in combinatorial problems, where annotations in the form of (near-) optimal information are generally produced with expensive (exact) solvers. To contrast labeling cost and improve generalization, Semi-Supervised [45] and Self-Supervised [31] are gaining popularity due to their ability to learn from unlabeled data. Despite this premise, little to no application of these paradigms can be found in the JSP and the combinatorial literature [5, 9].
Motivated by these observations, we introduce a novel self-supervised training strategy that is less complex than most RL algorithms, does not require the formulation of the Markov Decision Process, and relies only on model-generated information, thus removing the need for expensive annotations. Our proposal is based on two assumptions: i) we suppose to be able to generate multiple solutions for a problem, a common characteristic of generative models [37]; and ii) we suppose it is possible to discriminate solutions based on the problem objective. When these assumptions are met, we train a model by generating multiple solutions and using the best one according to the problem objective as a pseudo-label [30]. This procedure borrows from Semi-Supervised the idea of pseudo-labeling, but does not require any external and expansive annotation as in Self-Supervised learning. Hence, we refer to it as a Self-Labeling training strategy.
We demonstrate the viability of our training procedure in the context of the JSP. Similar to PDRs, we cast the generation of solutions as a sequence of decisions, where at each decision one operation is scheduled in the solution under construction. This is achieved with a popular architecture known as Pointer Network (PN) for dealing with sequences of decisions in combinatorial problems [47, 27, 49]. We train our PN on 30 000 instances of different sizes by generating multiple parallel solutions and using the one with the minimum makespan to update the model. This training strategy produces models outperforming state-of-the-art learning proposals and other conventional JSP algorithms on two popular benchmark sets. Additionally, our methodology proved to be robust and effective in various training regimes. Summarizing, the contributions of this work are:
  • We propose a novel Self-Labeling training strategy. Due to its assumptions, this effective strategy for the JSP can be applied as-is to other combinatorial problems.
  • We present a generative model based on the Pointer Network which produces high-quality JSP solutions in a matter of seconds when trained with our strategy.
The remainder of this work is organized as follows: Section 2 reviews related ML works; Section 3 formalizes the JSP; Section 4 introduces the proposed PN and training strategy; Section 5 compares our trained PN with other approaches; Section 6 studies aspects of our proposal; and Section 7 and 8 close with general considerations, some limitations, and potential future directions.
2 Related Works
ML approaches have been recently investigated to address issues of traditional JSP methodologies.
Motivated by the success of supervised learning, various studies have employed exact methods to generate optimality information for synthetic JSP instances, enabling the learning of valuable scheduling patterns. For example, [26] used imitation learning [5] to derive superior dispatching rules from optimal solutions, highlighting the limitations of learning solely from optimal solutions. [28] proposed a joint utilization of optimal solutions and Lagrangian terms during training to better capture JSP constraints. Whereas, in [14], a Recurrent Neural Network (RNN) learns to predict the quality of machine permutations, generated during training with a MIP, for guiding a metaheuristic. Despite their quality, these proposals are limited by their dependence on costly optimality information.
To avoid the need for optimality, other works have turned to reinforcement learning. The advantage of RL lies in its independence from costly optimal information, requiring only the correct formulation of the Markov Decision Process [42]. Several works focused on creating better neural PDRs, showcasing the effectiveness of policy-based methods [49, 38, 25, 11]. All these methods adopted a Proximal Policy Optimization algorithm [40] and proposed different architectures or training variations. For instance, [49, 38] adopted Graph Neural Networks (GNN), [11] proposed a rather complex Transformer [46], while [25] used RNNs and curriculum learning [53]. Differently, [20] presented a Double Deep Q-Network approach, proving the applicability of value-based methods.
Other works employed RL to enhance other algorithmic approaches. In [12], an iterative RL-based approach was proposed to rewrite regions of JSP solutions selected by a learned policy. Whereas [17] utilized a Deep Q-learning approach to control three points of interventions within metaheuristics, while [50] recently presented a RL-guided improvement heuristic. Differently, [44] presented a hybrid imitation learning and policy gradient approach coupled with Constraint Programming (CP) for outperforming PDRs and a CP solver.
3 Job Shop
In the Job Shop Problem, we are given a set of jobs $J=\{J_{1},\dots,J_{n}\}$, a set of machines $M=\{M_{1},\dots,M_{m}\}$, and a set of operations $O=\{1,\dots,o\}$. Each job $j\in J$ comprises a sequence of $m_{j}$ consecutive operations $O_{j}=(l_{j},\,\dots,l_{j}+m_{j})\subset O$ indicating the order in which the job must be performed on machines, where $l_{j}$ is the index of the first operation of the job ($l_{j}=1+\sum_{i=1}^{j-1}m_{i}$). An operation $i\in O$ has to be performed on machine $\mu_{i}\in M$ for an uninterrupted amount of time $\tau_{i}\in\mathbb{R}_{\geq 0}$. Additionally, machines can handle one operation at a time. The objective of the JSP is to provide an order to operations on each machine, such that the precedences within operations are respected and the time required to complete all jobs (makespan) is minimized. Formally, we indicate with $\pi$ a JSP solution comprising a permutation of operations for each machine and use $C_{i}(\pi)$ to identify the completion time of operation $i\in O$ in $\pi$.
A JSP instance can be represented with a disjunctive graph $G=(V,A,E)$ [3], where: $V$ contains one vertex for each operation $i\in O$; $A$ is the set of conjunctive (directed) arcs connecting consecutive operations of jobs (these arcs reflect the order in which operations must be performed); $E$ is the set of disjunctive (undirected) edges connecting operations to perform on the same machines. When the JSP is represented as a disjunctive graph $G$, finding a solution means providing a direction to all the edges in $E$, such that the resulting graph is directed and acyclic (all precedences are respected). As in related works [49, 38], we augment the disjunctive graph by associating to each vertex a set of 15 features $x_{i}\in\mathbb{R}^{15}$ describing information about operation $i\in O$. For lack of space, we present these features in Appendix A.
Figure 1: The sequences of decisions for constructing solutions in a JSP instance with two jobs ($J_{1}$ and $J_{2}$) and two machines (identified with different colors).
3.1 Constructing JSP Solutions
Feasible JSP solutions can be constructed step-by-step as a sequence of decisions, a common approach adopted by PDRs [39] and RL algorithms [49, 25]. The construction of solutions can be visualized using a branch-decision tree, shown at the bottom of Figure 1. Each path from the root node ($R$) to a leaf node presents a particular sequence of $o=|O|$ decisions and leads to a valid JSP solution, such as the one highlighted in gray. At every node in the tree, one uncompleted job needs to be selected, and its ready operation is scheduled in the partial solution being constructed. Due to the constraints imposed by the jobs (i.e., the conjunctive arcs) and the partial solution $\pi_{t}$ constructed up to decision $t$, there is at most one operation $o(t,\,j)\in O_{j}$ that is ready to be scheduled for any job $j\in J$. Thus, by selecting a job $j$, we uniquely identify an operation $o(t,\,j)$ that will be scheduled by appending it to the partial permutation of its machine. It is important to note that once a job is completed, it cannot be selected again, and such a situation is identified with a cross in Figure 1.
We also highlight that there exist multiple paths in the branch-decision tree that lead to the same solution. This indicates that, within the context of the JSP, certain solutions are more probable than others. In some instances, due to the distribution of processing times $\tau_{i}$, near-optimal solutions may be found among the less likely paths, making the construction of quality solutions a more complex task. This explains why it is more difficult for rigid constructive heuristics like standard PDRs to create high-quality solutions when such solutions result from very unlikely paths.
4 Pointer Network & Self-Labeling Strategy
Herein, we present the Pointer Network used to tackle the JSP and the proposed self-supervised training strategy.
4.1 Proposed Pointer Network
As described in Section 3.1, we tackle the JSP as a sequence of decisions. Thus, we propose a Pointer Network [47] whose goal is to select the right job at each decision step $t$. Recall that the PN is an encoder-decoder architecture to generate sequences of decisions in combinatorial problems [34, 27]. Formally, our PN learns a function $f_{\theta}(\cdot)$, parametrized by $\theta$, that estimates the probability $p_{\theta}(\pi|G)$ of a solution $\pi$ of being of high-quality as a product of probabilities:
$p_{\theta}(\pi|G)=\prod_{t=1}^{o}f_{\theta}(j\,|\,\pi_{t},G),$ (1)
where $f_{\theta}(j\,|\pi_{t},G)$ gives the probability of selecting job $j\in J$ conditioned on the instance $G$ and the partial solution $\pi_{t}$ at step $t$. By learning $f_{\theta}$, it is possible to construct solutions autoregressively with a high chance of being of high quality.
Our PN works by taking in input JSP instances represented as a disjunctive graph $G$, and produces single deterministic or multiple randomized solutions depending on the adopted construction strategy. The encoder of our PN creates with a single forward computation of $G$ an embedded representation for all the operations in $O$. Whereas the decoder uses the embeddings of operations and the partial solution $\pi_{t}$ to produce a probability for each job of being selected at step $t$. After selecting a job $j$, its ready operation $o(t,\,j)$ is scheduled in $\pi_{t}$ to produce the partial solution for step $t+1$. In the following, we present the details of both the encoder and decoder.
Encoder.
Its role is to encode meaningful patterns within an instance into the embedded representation of operations. The encoder can be embodied by any architecture, like a Feedforward Neural Network (FNN), that transforms the feature of operations $x_{i}\in\mathbb{R}^{15}$ (see Appendix A for the features) into embeddings $e_{i}\in\mathbb{R}^{h}$, without necessarily accounting for information contained in the graph $G$. As in related works [49, 38], we also encode the relationships among operations present in the disjunctive graph. Thus, we equip the encoder with Graph Neural Networks [48], which enables the embeddings to incorporate information related to the topological structure of $G$. In our encoder, we stack two layers of Graph Attention Network [8] (GAT) as follows:
$e_{i}=[x_{i}\>||\>\sigma(\text{GAT}_{\texttt{2}}(\,[x_{i}\;||\;\sigma(\text{GAT}_{\texttt{1}}(x_{i},\,G))],\,G)],$ (2)
where $\sigma$ is the ReLU non-linearity and ||||| | stands for the concatenation operation.
Table 1: The features of a context vector $c_{j}\in\mathbb{R}^{11}$ that describes the status of a job $j$ within a partial solution $\pi_{t}$. Recall that $o(t,j)$ is the ready operation of job $j$ at step $t$, $\mu_{o(t,\,j)}$ its machine, and $o(t,j)-1$ its predecessor.
ID Description Rationale
1 $C_{o(t,\,j)-1}(\pi_{t})$ minus the completion time of machine $\mu_{o(t,\,j)}$. The idle time of job $j$ if scheduled on its next machine $\mu_{o(t,\,j)}$ at time $t$. Negative if the job could have started earlier.
2 $C_{o(t,\,j)-1}(\pi_{t})$ divided by the makespan of $\pi_{t}$. How close job $j$ is to the makespan of the partial solution $\pi_{t}$.
3 $C_{o(t,\,j)-1}(\pi_{t})$ minus the average completion time of all jobs. How early/late job $j$ completes w.r.t. the average among all jobs in $\pi_{t}$.
4 - 6 The difference between $C_{o(t,\,j)-1}(\pi_{t})$ and the $1^{st}$, $2^{nd}$, and $3^{rd}$ quartile computed among the completion time of all jobs. Describe the relative completion of $j$ w.r.t. other jobs. These features complement that with ID = $3$.
7 The completion time of machine $\mu_{o(t,\,j)}$ divided by the makespan of the partial solution $\pi_{t}$. How close the completion time of machine $\mu_{o(t,\,j)}$ is to the makespan of the partial solution $\pi_{t}$.
8 The completion time of machine $\mu_{o(t,\,j)}$ minus the average completion of all machines in $\pi_{t}$. How early/late machine $\mu_{o(t,\,j)}$ completes w.r.t. the average among machines in $\pi_{t}$.
9 - 11 The difference between the completion of $\mu_{o(t,\,j)}$ and the $1^{st}$, $2^{nd}$, and $3^{rd}$ quartile among the completion time of all machines. Describe the relative completion time of machine $\mu_{o(t,\,j)}$ w.r.t. all the other machines. These features complement that with ID = $8$.
Decoder.
Its role is to produce at any step $t$ the probability of selecting each job with two logically distinct components:
  • Memory Network: generates a state $s_{j}\in\mathbb{R}^{d}$ for each job $j\in J$ from the partial solution $\pi_{t}$. This is achieved by first extracting from $\pi_{t}$ a context vector $c_{j}$ (similarly to [27]), which contains eleven hand-crafted features providing useful cues about job $j$ in the partial solution. Note that these features are an integral part of our architecture and we refer to Table 1 for their definition and meaning. Then, these vectors are fed into a Multi-Head Attention layer (MHA) [46] followed by a non-linear projection to produce jobs’ states:
    $s_{j}=\text{ReLU}([c_{j}\,W_{\texttt{1}}+\underset{b\in J}{\text{MHA}}(c_{b}\,W_{\texttt{1}})]\,W_{\texttt{2}}),$ (3)
    where $W_{\texttt{1}}$ and $W_{\texttt{2}}$ are projection matrices. Note that we use the MHA to consider the context of all jobs when producing the state for a specific one, similarly to [11].
  • Classifier Network: outputs the probability $p_{j}$ of selecting a job $j$ by combining the embedding $e_{o(t,\,j)}$ of its ready operation and the state $s_{j}$ produced by the memory network. To achieve this, we first concatenate the embeddings $e_{o(t,\,j)}$ with the states $s_{j}$ and apply an FNN:
    $z_{j}=\text{FNN}([e_{o(t,\,j)}\,||\,s_{j}]).$ (4)
    Then, these scores $z_{j}\in\mathbb{R}$ are transformed into probabilities with a Softmax function: $p_{j}=e^{z_{j}}\,/\,\sum\nolimits_{b\in J}e^{z_{b}}$. Finally, the decision of which job to select at $t$ is made with a sampling strategy, as explained next.
Sampling solutions.
To generate solutions, we employ a probabilistic approach for deciding the job selected at any step. Specifically, we randomly sample a job $j$ with a probability proportional to $p_{j}$, which is produced at step $t$ by our decoder. We also prevent the selection of completed jobs by setting their probabilities to zero before sampling, i.e., we set $p_{j}=0$ for completed $j$. Note how this probabilistic selection is autoregressive, meaning that sampling a job depends on $p_{j}$ which is a function of $e_{o(t,\,j)}$ and $s_{j}$ resulting from earlier decisions (see Eq. 3 and 4). There exist various strategies for sampling from autoregressive models, including top-k [23], nucleus [23], and random sampling [4]. After preliminary analysis, we found that these strategies exhibit similar effectiveness in our setting. Therefore, we adopt the simplest strategy for training and testing, which is random sampling, i.e., sampling a job $j$ with probability $p_{j}$ at random. Lastly, we use our PN to generate $\beta$ parallel solutions for an instance. This is achieved by keeping $\beta$ separate partial solutions and sampling at any step a job for each one independently.
4.2 Self-Labeling Training Strategy
We propose a simple self-supervised training strategy that resorts only to the model being trained and does not require optimality information nor the formulation of the Markov Decision Process. Our strategy exploits two aspects: the ability of generative models to construct multiple (parallel) solutions, and the possibility of discriminating solutions based on their objective values, such as the makespan in the JSP. With these two ingredients, we design a procedure that at each iteration generates multiple solutions and uses the best one as a pseudo-label [30].
More in detail, for each training instance $G$, we generate with the PN a set of $\beta$ different solutions. We do that by keeping $\beta$ partial solutions in parallel and independently sample for each one the next job from the probabilities generated by the decoder. Once the solutions have been generated, we take the one with the minimum makespan $\overline{\pi}$ and use it to compute the Self-Labeling loss ($\mathcal{L}_{\operatorname{SL}}$) by minimizing the cross-entropy on all the decision steps as follows:
$\mathcal{L}_{\operatorname{SL}}(\overline{\pi})=-\frac{1}{o}\sum_{t=1}^{o}\log f_{\theta}(y_{t}\,|\,\pi_{t},G),$ (5)
where $o$ is the number of operations in the instance and $y_{t}\in J$ is the index of the job selected at the decision step $t$ while constructing the solution $\overline{\pi}$.
The rationale of this training schema is to collect knowledge about sequences of decisions leading to high-quality solutions and save it in the model parameters $\theta$. This is done by selecting the best-generated solution $\overline{\pi}$ for an instance $G$ and maximizing its likelihood with Eq. 5 as in supervised learning, i.e., by minimizing cross-entropy. Through iterative exposure to different training instances, the model refines its ability to effectively solve the JSP.
Although our strategy was developed independently, attentive readers may find a resemblance with the Cross-Entropy Method (CEM), a stochastic and derivative-free optimization method [15]. Typically applied to optimize parametric models such as Bernoulli and Gaussian mixtures models [52], the CEM relies on a maximum likelihood approach to either estimate a random variable or optimize the objective function of a problem [7]. According to Algorithm 2.2 in [7], the CEM independently samples $N=\beta$ solutions, selects a subset based on the problem’s objective ($\hat{\gamma}=S_{(N)}$ in our case), and updates the model’s parameters.
Unlike the CEM, which independently tackles different instances by optimizing a separate model for each, our strategy trains a single model on many instances for learning how to globally solve the JSP. Moreover, we adopt a more complex parametric model rather than mixture models, always select a single solution for updating, and resorting to the gradient descent for updating parameters. Notably, our strategy also differs from CEM applications to model-based RL, e.g., [52], as we do not rely on rewards in Eq. 5. Summarizing, our strategy shares the idea of sampling and selecting with the CEM, but operates globally as a fully supervised learning once target solutions are identified. Hence, we named it: self-labeling training strategy.
Other self-labeling approaches can be found e.g. in [10, 2], where the former uses K-Means to generate pseudo-labels, and the latter assigns labels to equally partition data with an optimal transportation problem. As highlighted in [2], these approaches may incur in degenerative solutions, i.e., solutions that trivially minimize Eq. 5, such as producing the same solution despite the input instance in our case. We remark that our strategy avoids such solutions by jointly using a probabilistic generation process and the objective value (makespan) of solutions to select the best pseudo-label $\overline{\pi}$.
5 Results
This section describes implementation details as well as selected competitors, and then presents the results. All the experiments were performed on an Ubuntu 22.04 machine equipped with an Intel Core i9-11900K and an NVIDIA GeForce RTX 3090 having 24GB of memory.
5.1 Experimental Setup
Dataset & Benchmarks.
[…]
Architecture.
In all our experiments, we configure the PN of Section 4.1 in the same way. Our encoder consists of two GAT layers [8], both with 3 attention heads and Leaky slope at 0.15. In $\text{GAT}_{\texttt{1}}$, we set the size of each head to $64$ and concatenate their outputs; while in $\text{GAT}_{\texttt{2}}$, we increase the head’s size to $128$ and average their output to produce $e_{i}\in\mathbb{R}^{143}$ ($h=15+128$). Inside the decoder’s memory network, the MHA layer follows [46] except it concatenates the output of 3 heads with 64 neurons each, while $W_{\texttt{1}}\in\mathbb{R}^{11\times 192}$ and $W_{\texttt{2}}\in\mathbb{R}^{192\times 128}$ use 192 and 128 neurons, respectively. Thus, in Eq. 3, states $s_{j}\in\mathbb{R}^{d}$ have size $d=128$. Regarding the classifier, the FNN features a dense layer with 128 neurons activated through the Leaky-ReLU (slope $=0.15$) and a final linear with 1 neuron.
Training.
We train the PN with the strategy outlined in Section 4.2 on our dataset for $20$ epochs with the Adam optimizer [19] and learning rate at $0.0002$. During training, we save the parameters producing the lower average makespan on a hold-out set comprising 100 random instances of each shape. Additionally, we fix the number of sampled solutions $\beta$ in training to $256$, and use a batch size of 16, hence accumulating the gradient for 16 instances before updating $\theta$. Training this way roughly takes 120 hours, around 6 hours per epoch.
Competitors.
We compare the effectiveness of our methodology against various types of conventional and ML competitors reviewed in Section 2. We divide competitors as follows:
  • Greedy constructives generate a single solution for any input instance. We consider two cornerstone ML works for the JSP: the actor-critic (L2D) of [49] and the recent curriculum approach (CL) of [25]. We also coded the INSertion Algorithm (INSA) of [35] and three standard dispatching rules that prioritize jobs based on the Shortest Processing Time (SPT); Most Work Remaining (MWR); and Most Operation Remaining (MOR).
  • Multiple (Randomized) constructives generate many solutions for an instance by introducing a controlled randomization in the selection process, a simple strategy for enhancing constructive algorithms [47, 25]. We consider the randomized results of CL and those of L2D. Since only greedy results are disclosed for L2D, we used the open-source code and trained model to generate randomized solutions as described at the bottom of Section 4.1. We also consider the three dispatching rules above, randomized by arbitrarily selecting the operation to schedule among the three with higher priority, and the Shifting Bottleneck Heuristic (SBH) [1] that creates multiple solutions while optimizing. All these approaches were seeded with 12345, generate $\beta=128$ solutions, and return the one with minimum makespan.
  • Non-constructive approaches do not rely on a pure constructive strategy for creating JSP solutions. We include two RL-enhanced metaheuristics: the NLSA in [18] (200 iterations) and the recent L2S proposal of [50] visiting 500 (L2S500) and 5000 (L2S5k) solutions. We also consider two state-of-the-art solvers for the JSP: Gurobi 9.5 (MIP) solving the disjunctive formulation of [29] and the CP-Sat (CP) of Google OR tools 9.8, both executing with a time limit of 3600 seconds.
For lack of space, we compare our PN against other learning proposals in Appendix B.
5.2 Performance on Benchmarks
Table 2: The average PGs of the algorithms on the two benchmark sets. For each row, we highlight in blue and bold the lowest (best) constructive and non-constructive gap, respectively. Shapes marked with * are larger than those seen in training by our SPN.
Greedy Constructives Multiple (Randomized) Constructives Non-constructive
Shape
Taillard’s benchmark $15$$\times$ 1515 × 15
$\ 20\times 15$
$20$$\times$ 2020 × 20
$\ 30\times 15^{*}$
$\ 30\times 20^{*}$
$\ 50\times 15^{*}$
$\ 50\times 20^{*}$
$100\times 20^{*}$
Avg
Demirkol’s benchmark $20\times 15$
$20\times 20$
$30\times 15^{*}$
$30\times 20^{*}$
$40\times 15^{*}$
$40\times 20^{*}$
$50\times 15^{*}$
$50\times 20^{*}$
Avg
PDRs RL Our
SPT MWR MOR INSA L2D CL SPN
26.09 19.09 20.40 14.43 25.96 14.26 13.77
32.30 23.33 24.86 18.88 30.03 16.52 14.96
28.27 21.81 22.89 17.29 31.61 17.27 15.17
35.04 24.07 22.89 21.14 33.00 18.52 17.09
33.45 24.82 26.75 22.64 33.62 21.47 18.49
24.01 16.45 17.59 15.89 22.38 12.23 10.06
25.59 17.84 16.79 20.34 26.51 13.24 11.55
14.05 8.33 8.71 13.49 13.61 5.86 5.85
27.35 19.47 20.10 18.01 27.09 14.92 13.37
27.94 27.54 30.68 24.18 39.03 - 17.99
33.19 26.80 26.26 21.27 37.73 - 19.36
31.15 31.89 36.86 26.51 42.03 - 21.75
34.37 32.06 32.34 28.50 39.69 - 25.73
25.26 27.00 35.85 24.68 35.55 - 17.51
33.81 32.28 35.89 29.44 39.64 - 22.19
24.30 27.65 34.88 23.05 36.46 - 15.67
30.01 30.34 36.76 30.22 39.52 - 22.44
30.00 29.45 33.69 25.98 38.71 - 20.33
PDRs ($\beta=128$) RL ($\beta=128$) Our
SPT MWR MOR SBH L2D CL SPN128 SPN512
13.47 13.47 12.53 11.56 17.11 9.02 7.24 6.52
18.45 17.17 16.38 10.94 23.74 10.58 9.31 8.81
16.74 15.59 14.69 13.64 22.61 10.87 9.95 9.04
23.23 18.96 17.30 15.09 24.35 13.98 10.97 10.61
23.59 19.94 20.44 17.74 28.40 16.09 13.43 12.66
14.14 13.51 13.52 13.16 17.06 9.32 5.47 4.93
17.65 14.57 13.96 20.79 20.44 9.89 8.35 7.58
10.39 7.02 7.11 15.57 13.30 3.96 2.32 2.06
17.21 15.03 14.49 14.81 20.50 10.46 8.38 7.78
17.17 22.28 23.85 12.51 29.27 19.43 12.00 11.32
18.83 18.89 21.61 13.52 27.10 15.97 13.51 12.33
20.73 26.79 30.73 18.15 33.99 16.48 14.41 14.01
23.31 26.12 28.34 17.31 33.59 20.18 17.10 15.80
17.90 23.15 30.25 14.43 31.54 17.62 11.71 10.91
24.46 27.58 31.85 20.28 35.80 25.64 15.99 14.81
17.70 24.07 30.97 18.23 32.73 21.74 11.22 10.64
23.50 26.75 32.71 22.75 36.12 15.17 15.81 15.00
20.44 24.46 28.79 17.15 32.52 19.03 13.97 13.10
Approaches
L2S500 NLSA L2S5k MIP CP
9.3 7.74 6.2 0.07 0.1
11.6 12.16 8.3 3.22 0.2
12.4 11.54 9.0 2.88 0.7
14.7 14.13 9.0 10.66 2.1
17.5 16.35 12.6 13.18 2.8
11.0 11.01 4.6 12.25 0.0
13.0 11.25 6.5 13.59 2.8
7.9 5.90 3.0 10.97 3.9
12.2 11.26 7.40 8.36 1.6
- - - 5.29 1.8
- - - 4.69 1.9
- - - 14.16 2.5
- - - 16.67 4.4
- - - 16.33 4.1
- - - 22.50 4.6
- - - 14.90 3.8
- - - 22.49 4.8
- - - 14.63 3.5
This section evaluates the performance of our Self-labeling Pointer Network (SPN), configured and trained as explained in Section 5.1. When contrasted with greedy constructive approaches, our SPN generates a single solution by picking the job with the highest probability. In the other comparisons, we randomly sample 128 (SPN128) and 512 (SPN512) solutions as explained in Section 4.1.
We coded in Python 3.9 and PyTorch 13.1 our SPN, dispatching rules, INSA, SBH, MIP, and CP. While we report results from original papers of all the ML proposals but L2D, for which we used the open-source code for generating randomized solutions. Lastly, we evaluate performance on each benchmark instance in terms of the Percentage Gap: $\operatorname{PG}=100\cdot(C_{alg}\,/\,C_{ub}-1)$, where $C_{alg}$ is the makespan produced by an algorithm and $C_{ub}$ is either the instance optimal or best-known makespan.
Table 2 presents the comparison of algorithms on Taillard’s and Demirkol’s benchmarks, each arranged in a distinct horizontal section. The table is vertically divided into Greedy Constructive, Multiple (Randomized) Constructive, and Non-constructive approaches; with the results of algorithms categorized accordingly. Each row reports the average PG on a specific instance shape while the last row (Avg) reports the average gap across all instances, regardless of their shapes.
Table 2 shows that our SPN and CL produce lower gaps than PDRs, INSA, and SBH in both the greedy and randomized case, demonstrating the superiority of neural constructive approaches. Surprisingly, INSA and PDRs outperform L2D. This is likely related to how PDRs were coded in [49], where ours align with those in [25]. Focusing on our SPN, we observe that it consistently archives lower gaps than all the greedy approaches and, when applied in a randomized manner (SPN128), it outperforms all the randomized constructives. Notably, the SPN generalizes well to instance shapes larger than training ones (marked with * in Table 2), as its gaps do not progressively increase on such shapes.
Our SPN is also competitive when compared with non-constructive approaches. The SPN512 largely outperforms the NLSA and L2S500 metaheuristics, and roughly align with L2S5k visiting 5000 solutions. On shapes larger than $20\times 20$, the SPN512 remarkably achieves lower gaps with just a few seconds of computations than a MIP executing for 3600 sec. Therefore, we conclude that our proposed SPN and training strategy are effective in solving the JSP.
As already observed in [25, 50], our SPN and other neural constructive approaches are generally outperformed by CP and state-of-the-art metaheuristics, e.g., [35]. Although the performance gap is steadily narrowing, the higher complexity of CP solvers renders them more powerful, albeit at the cost of longer execution time (often dozens of minutes). Therefore, neural constructive approaches may be preferable whenever a quality solution must be provided in a few minutes.
We refer the reader to Appendix D and Appendix E for statistical considerations and a comparison between our SPN and CP on very large instances, respectively.
6 Ablations
This section studies additional aspects of our proposed methodology. For these evaluations, we adopt the same setting described in Section 5.1 and we explicitly indicate the varied parameters.
RL Self-Labeling
[…]
Taillard 17.30 10.46 10.70 9.51 8.38
Demirkol - 19.01 20.87 14.91 13.97
Table 3: The average gaps when sampling 128 solutions from architectures trained without and with self-labeling. ${}_{\text{UCL}}$ is the model obtained in [25] by training with reward-to-go on random instance shapes (no curriculum learning) and CL is similarly obtained by applying curriculum learning.
Figure 2: SPN validation curves when trained with PPO and Self-Labeling in the same training setting of [49].
6.1 Self-labeling other Architectures
[…]
6.2 Comparison with Proximal Policy Optimization
Section 6.1 proved the quality of the SPN when trained with our self-labeling strategy. As Proximal Policy Optimization (PPO) is extensively used to train neural algorithms for scheduling problems (see Section 2), we contrast the SPN’s performance when trained with PPO and our Self-Labeling strategy (SL-$\beta$ with $\beta\in\{32,64,128\}$). For this evaluation, we adopted the same hyper-parameters of [49] for both PPO and SL by training on 40000 random instances of shape 30x20 (same training of L2D). We only double the batch size to 8 for making the training less noisy. Figure 2 reports the validation curves obtained by testing the SPN on Taillard’s benchmark every 100 training steps. We test by sampling 128 solutions and we also include the randomized results of L2D (dashed line) as a baseline for comparisons. From Figure 2, we observe that training with our self-labeling strategy results in faster convergence and produces better final models.
We justify this by hypothesizing that the reward received from partial solutions (makespan increments) may not always provide reliable guidance on how to construct the best final solution. While constructing a solution, the critical path may change with implications on past rewards. Our self-labeling strategy does not rely on partial rewards and may avoid such a source of additional “noise". Note that this is only an intuition, proving it would require a deeper analysis, which we refer to in future works. Finally, we stress that we do not claim our self-labeling strategy is superior to PPO. Instead, we believe it offers an alternative that provides a fresh perspective and potential for integration with existing RL algorithms to advance the neural combinatorial optimization field.
Figure 3: The performance of the SPN when trained by sampling $\beta$ solutions. For each shape, we report the average PG on instances of both benchmarks by sampling $512$ solutions during testing.
Figure 4: The performance of the SPN for varying numbers of sampled solutions $\beta$ at test time. For each shape, we report the average PG on instances of both benchmarks.
6.3 The effect of $\beta$ on Training
As our self-labeling strategy is based on sampling, we assess the impact of sampling a different number of solutions $\beta$ while training. We retrain a new SPN as described in Section 5.1 with a number of solutions $\beta\in\{32,64,128,256\}$, where we stop at 256 as the memory usage with larger values becomes impractical. We test the resulting models by sampling 512 solutions on all the instances of both benchmarks for a broader assessment. Figure 3 reports the average PG (the lower the better) of the trained SPN (colored markers) on each shape. To ease comparisons, we also include the results of CL (dashed line), the second-best constructive proposal in Table 2, and the MIP (dotted line).
Overall, we see that training by sampling more solutions slightly improves the model’s overall performance, as outlined by lower PGs for increasing $\beta$. We also observe that such improvement is less marked on shapes seen in training, such as in $15\times 15,20\times 15$, and $20\times 20$ shapes, while is more marked on others. This suggests that training by sampling more solutions results in better generalization, although it is more memory-demanding. However, the trends observed for the SPN in Table 2 remain consistent with smaller $\beta$, demonstrating the robustness of our self-labeling strategy.
6.4 The effect of $\beta$ on Testing
We also assess how the number of sampled solutions $\beta$ impacts the performance of our SPN at test time. To evaluate such an impact, we plot in Figure 4 the average PG on different shapes for varying $\beta\in\{32,64,128,256,512\}$. For this analysis, we use the SPN trained by sampling 256 solutions, the same used in Table 2. As in Section 6.3, we also report the results of CL and MIP to ease the comparison.
Despite the reduced number of solutions, the SPN remains a better alternative than CL, one of the best RL proposals, and still outperforms the MIP on medium and large instances. Not surprisingly, by sampling more solutions the performance of the SPN improves, but this also increases the model execution times. Although we verified that sampling more than 512 solutions further improves results, we decided to stop at $\beta=512$ as a good trade-off between performance and time. We refer the reader to Appendix F for other timing considerations.
7 Limitations
Despite the proven effectiveness of our self-labeling strategy, it is important to note that only one of the sampled solutions per training instance is used to update the model’s parameters. This approach may be sub-optimal from an efficiency standpoint. Therefore, we see significant potential in hybridizing our self-labeling strategy with existing (RL) methods to mitigate this inefficiency. Moreover, sampling multiple solutions during training requires substantial memory, which can limit batch sizes. Although we have shown that our self-labeling strategy remains effective with a small number of sampled solutions (see Section 6.3), increasing the number of samples accelerates training convergence and improves the quality of the resulting model. Consequently, developing new sampling strategies that can construct high-quality solutions without generating numerous random ones is a promising future direction. Such advancements could reduce memory usage and further enhance our methodology.
8 Conclusions
This work outlines a novel self-labeling strategy to train generative models for the Job Shop and potentially other combinatorial problems. Our contribution is twofold: an efficient model capable of generating multiple parallel solutions, and the training strategy that solely relies on model-generated solutions. Despite its simplicity, our methodology significantly outperforms PDRs and many learning proposals, even surpassing a powerful MIP solver. However, due to its constructing nature, it still lags behind state-of-the-art methods like constraint programming solvers, which nevertheless require more time. We also proved the resiliency of our approach to different parameters and architectures.
More broadly, while the right training strategy and architecture are crucial, we emphasize the paramount importance of providing the model with the right features and state representation to create effective neural algorithms for combinatorial optimization problems. A common trend in the literature is to let the model learn its own representation by applying it to the limited information contained in the problem instance. However, our experience suggests that incorporating additional problem-specific information, like the features described in Table 1 for the JSP, can significantly enhance performance, often more than adopting variations in the training strategies.
In the future, we intend to apply our methodology to other combinatorial problems and explore combinations of our self-supervised strategy with existing RL approaches.
Appendix A Disjunctive Graph Features
Following ML works reviewed in Section 2, we augment the standard disjunctive graph representation of a JSP instance by incorporating additional features for each vertex $i\in V$. Table 4 details these features. Some features, such as features 1-3, have been adapted from previous works [49, 38], while others have been introduced by us to better model JSP concepts like machines and jobs. For instance, features 4-6 take the same value for operations of the same job, emphasizing which operations belong to different jobs, while features 7-9 do the same for operations to be performed on the same machine.
Table 4: The features $x_{i}\in\mathbb{R}^{15}$ associated with a vertex $i\in V$ of the disjunctive graph $G$ provide information about operation $i$ in the instance. Recall that $\tau_{i}$ represents the processing time of operation $i$, and $\mu_{i}$ denotes the machine on which the operation is performed.
ID Description
1 The processing time $\tau_{i}$ of the operation.
2 The completion of job $j$ up to $i$: $\sum_{b=l_{j}}^{i}\tau_{b}\,/\,\sum_{b\in O_{j}}\tau_{b}$.
3 The remainder of job $j$ after $i$: $\sum_{b=i+1}^{l_{j}+m_{j}}\tau_{b}\,/\,\sum_{b\in O_{j}}\tau_{b}$.
4-6 The $1^{st}$, $2^{nd}$, and $3^{rd}$ quartile among processing times of operations on job $j$.
7-9 The $1^{st}$, $2^{nd}$, and $3^{rd}$ quartile among processing times of operations on machine $\mu_{i}$.
10-12 The difference between $\tau_{i}$ and the $1^{st}$, $2^{nd}$, and $3^{rd}$ quartile among processing times of operations in job $j$.
13-15 The difference between $\tau_{i}$ and the $1^{st}$, $2^{nd}$, and $3^{rd}$ quartile among processing times of operations on machine $\mu_{i}$.
Appendix B Additional Comparisons with Neural Algorithms
We additionally consider learning proposals reviewed in Section 2, which were tested only on a few benchmark instances or which could not be included in Table 2 due to how results are disclosed.
Regarding proposals not included in Table 2, we highlight that our SPN outperforms the Graph Neural Network in [38] which exhibits a significantly higher average gap of 19.50% on Taillard’s benchmark. Note that only qualitative results are disclosed in [38], we did our best to report them. Similarly, we remark that the SPN’s performance is also superior to the proposal in [44] (hCP), mixing RL and constraint programming. hCP achieves an average makespan of 2670 and 5701 on Taillard and Demirkol benchmarks respectively, while our greedy SPN yet obtains an average of 2642 and 5581.
As some other learning proposals like [20, 11] were tested only on certain benchmark instances, we provide an additional evaluation on such instances for a fair comparison. We include in Table 5 the average PGs of greedy constructive approaches on the two Taillard’s instances of each shape used in the Transformer (TRL) proposal of [11] and the Deep Q-Network (DQN) of [20].
Table 5: The average PG of greedy constructive approaches on the same two Taillard’s instances of the ten available for each shape, as utilized in [20, 11]. We highlight in bold the best PG on each instance shape.
Shape
$15$$\times$ 1515 × 15
$\ 20\times 15$
$20$$\times$ 2020 × 20
$\ 30\times 15^{*}$
$\ 30\times 20^{*}$
$\ 50\times 15^{*}$
$\ 50\times 20^{*}$
$100\times 20^{*}$
Avg
PDRs RL
SPT MWR MOR INSA TRL L2D CL DQN SPN
17.50 18.76 15.21 15.45 35.37 22.34 15.16 7.11 15.14
29.69 24.52 26.83 15.25 32.10 26.83 17.85 13.87 13.99
27.76 22.05 23.10 16.70 28.29 30.27 16.16 17.88 17.37
39.45 22.79 23.92 21.13 36.41 31.47 20.29 16.13 17.06
30.03 27.70 27.51 19.17 34.66 38.26 22.15 21.78 21.13
30.14 23.78 26.16 12.42 31.85 24.58 15.61 17.69 11.82
24.26 18.37 18.62 22.96 28.03 28.62 14.19 19.50 14.42
14.23 8.21 8.54 15.19 17.99 12.47 5.50 9.52 4.96
26.63 20.77 21.24 17.28 30.59 26.85 15.86 15.44 14.49
On this subset of instances, we see that TRL is the worst-performing approach and that DQN roughly aligns with CL. As our SPN is always the best or second best algorithm on each shape, it achieves the lowest overall average gap (Avg).
Appendix C Performance on Lawrence’s Benchmark
Lawrence’s benchmark has been extensively used in the JSP literature to develop various resolution approaches, e.g. [1, 35, 50]. This benchmark includes 40 instances of 8 different shapes and features smaller (and easier) instances compared to the Taillard’s and Demirkol’s ones. To complement the evaluation of Section 5.2 on small instances, we provide in Table 6 a comparison of approaches included in Table 2 on Lawrence’s benchmark. Note that we could not include CL and NLSA as they were not tested on these instances.
Table 6: The average PGs of coded algorithms on Lawrence’s benchmark. For each row, we highlight in blue and bold the lowest (best) constructive and non-constructive gap, respectively.
Greedy Constructives Multiple (Randomized) Constructives Non-constructive
Shape
Lawrence’s benchmark $10\times 5$
$15\times 5$
$20\times 5$
$10\times 10$
$15\times 10$
$20\times 10$
$30\times 10$
$15\times 15$
Avg
PDRs RL Our
SPT MWR MOR INSA L2D SPN
14.81 15.91 14.60 7.33 14.26 9.33
14.86 5.49 5.00 2.18 7.83 2.58
16.29 5.17 6.71 2.66 6.28 2.12
15.67 12.20 11.95 7.92 23.74 8.88
28.69 17.83 23.38 12.07 27.20 11.61
36.87 17.91 21.87 15.95 24.56 12.14
16.62 9.21 7.72 3.89 8.39 2.01
25.84 18.21 18.65 14.84 27.13 13.58
21.21 12.74 13.73 8.35 17.42 7.78
PDRs ($\beta=128$) RL Our
SPT MWR MOR SBH L2D SPN128 SPN512
3.78 9.02 5.64 1.96 8.80 1.86 1.07
2.26 1.19 1.10 0.00 2.77 0.00 0.00
5.84 1.43 2.60 0.07 3.08 0.00 0.00
8.54 7.56 6.85 5.69 10.42 3.12 2.51
15.24 11.05 11.64 7.56 16.18 5.22 5.02
17.67 12.54 12.87 5.82 18.34 6.85 5.59
8.52 3.47 2.25 0.00 6.76 0.00 0.00
12.33 12.17 12.84 8.44 17.42 6.77 5.59
9.27 7.30 6.96 3.69 10.47 2.98 2.47
Approaches
L2S500 L2S5k MIP CP
2.1 1.8 0.0 0.0
0.0 0.0 0.0 0.0
0.0 0.0 0.0 0.0
4.4 0.9 0.0 0.0
6.4 3.4 0.0 0.0
7.0 2.6 0.0 0.0
0.2 0.0 0.0 0.0
7.3 5.9 0.0 0.0
3.43 1.83 0.0 0.0
Table 6 shows that our SPN remains the best constructive approach for small instances. However, we observe that L2S5k achieves slightly better results than SPN512 at the cost of significantly longer execution times. Specifically, L2S5k always requires more than 70 sec on such instance shapes (as reported in [50]), whereas our SPN512 always completes in less than 0.5 sec (see also Appendix. F for timing considerations). Finally, we highlight that on small instances like those in Lawrence’s benchmark, exact methods such as MIP and CP can effectively solve them to optimality.
Appendix D Statistical Analysis
Table 7: The average PG and its standard deviation ($avg\,\pm\,std$) of the best four multiple (randomized) constructive and non-constructive algorithms on Taillard’s benchmark. We highlight in blue (bold) the best constructive (non-constructive) gap.
Shape
Taillard’s benchmark $15$$\times$ 1515 × 15
$\ 20\times 15$
$20$$\times$ 2020 × 20
$\ 30\times 15$
$\ 30\times 20$
$\ 50\times 15$
$\ 50\times 20$
$100\times 20$
Tot Avg $\pm$ Std
Multiple (Randomized) Constructive
MOR CL SPN128 SPN512
12.53± $\,\pm\,2.6$  9.02± $\,\pm\,2.0$  7.24± $\,\pm\,1.6$  6.52± $\,\pm\,1.3$
16.38± $\,\pm\,1.5$ 10.58± $\,\pm\,1.6$  9.31± $\,\pm\,1.8$  8.81± $\,\pm\,1.4$
14.69± $\,\pm\,1.7$ 10.87± $\,\pm\,1.6$  9.95± $\,\pm\,1.3$  9.04± $\,\pm\,1.0$
17.30± $\,\pm\,4.5$ 13.98± $\,\pm\,3.8$ 10.97± $\,\pm\,3.8$ 10.61± $\,\pm\,3.6$
20.44± $\,\pm\,2.2$ 16.09± $\,\pm\,2.0$ 13.43± $\,\pm\,1.6$ 12.66± $\,\pm\,1.4$
13.52± $\,\pm\,4.5$  9.32± $\,\pm\,3.8$  5.47± $\,\pm\,3.0$  4.93± $\,\pm\,2.7$
13.96± $\,\pm\,1.5$  9.89± $\,\pm\,2.2$  8.35± $\,\pm\,2.0$  7.58± $\,\pm\,2.0$
 7.11± $\,\pm\,2.9$  3.96± $\,\pm\,2.2$  2.32± $\,\pm\,1.1$  2.06± $\,\pm\,1.2$
14.49± $\,\pm\,4.6$ 10.46± $\,\pm\,4.2$ 8.38± $\,\pm\,3.8$ 7.78± $\,\pm\,3.6$
Non-Constructive Approaches
NLSA L2S5k MIP CP
 7.74± $\,\pm\,3.0$  6.2± $\,\pm\,1.4$  0.07± $\,\pm\,0.2$ 0.1± $\,\pm\,0.1$
12.16± $\,\pm\,1.7$  8.3± $\,\pm\,1.7$  3.22± $\,\pm\,1.9$ 0.2± $\,\pm\,0.4$
11.54± $\,\pm\,1.5$  9.0± $\,\pm\,2.2$  2.88± $\,\pm\,1.6$ 0.7± $\,\pm\,0.7$
14.13± $\,\pm\,3.5$  9.0± $\,\pm\,3.0$ 10.66± $\,\pm\,3.1$ 2.1± $\,\pm\,2.2$
16.35± $\,\pm\,2.4$ 12.6± $\,\pm\,2.1$ 13.18± $\,\pm\,2.9$ 2.8± $\,\pm\,1.5$
11.01± $\,\pm\,4.3$  4.6± $\,\pm\,3.1$ 12.25± $\,\pm\,3.9$ 0.0± $\,\pm\,0.0$
11.25± $\,\pm\,2.1$  6.5± $\,\pm\,1.9$ 13.59± $\,\pm\,4.1$ 2.8± $\,\pm\,1.9$
 5.90± $\,\pm\,2.3$  3.0± $\,\pm\,1.7$ 10.97± $\,\pm\,3.5$ 3.9± $\,\pm\,1.5$
11.26± $\,\pm\,4.1$ 7.40± $\,\pm\,3.7$ 8.36± $\,\pm\,5.9$ 1.6± $\,\pm\,1.9$
Due to space constraints, we evaluate algorithms in Table 2 solely in terms of average gaps. This approach is justified as the average trends are consistent with other statistical measures. For example, Table 7 reports the average and standard deviation of the best four multiple (randomized) constructive and non-constructive approaches on Taillard’s benchmark. We omit Demirkol’s instances since not all algorithms report results for them. Each row in Table 7 presents the average gap and standard deviation for instances of a specific shape, but the last row, which reports the average and standard deviation on all instances, regardless of their shapes. This table confirms that the observations made for Table 2 remain valid when considering standard deviations. Specifically, SPN512 remains the best constructive approach, aligning closely with the RL-enhanced metaheuristic L2S5k. The MIP still has lower performance than SPN512 while CP produces the best average gaps and standard deviations.
Appendix E Extremely Large Instances
Table 8: Average makespan ($avg\,\pm\,std$) and execution time of CP-Sat and our SPN (when sampling $\beta=512$ solutions) on very large instances.
Instance shape
200$\times$20 200$\times$40 500$\times$20 500$\times$40
CP-Sat $C_{max}$ 21848.5± $\,\pm\,331$ 23154.6± $\,\pm\,350$ 53572.7± $\,\pm\,743$ 546064.0± $\,\pm\,443$
$time$ 3600 $sec$ 3600 $sec$ 3600 $sec$ 3600 $sec$
SPN512 $C_{max}$ 21794.5± $\,\pm\,324$ 22795.0± $\,\pm\,277$ 53288.4± $\,\pm\,737$ 53942.9± $\,\pm\,425$
$time$ 31 $sec$ 64 $sec$ 164 $sec$ 343 $sec$
Gap sum from CP-Sat -3.30 -30.88 -6.30 -26.14
As shown in Table 2, constructive approaches are generally inferior to state-of-the-art metaheuristics [35] (see also [50]), Mixed Integer Programming (MIP) [29], and Constraint Programming (CP) solvers. These resolution methods adopt more complex frameworks that make them more powerful for solving combinatorial problems. However, neural constructive approaches are continuously closing the gap as remarked by the SPN’s performance matching that of MIP solvers. Herein, we also prove that our SPN generally produces better results than CP solvers on very large instances.
To this end, we randomly generated 20 JSP instances with shapes $\{200\times 20,200\times 40,500\times 20,500\times 40\}$ and solved them with the CP-Sat solver of Google OR-tools 9.8, executing for 1 hour, and our SPN, sampling $\beta=512$ solutions. Table 8 reports the average execution time ($time$), the average makespan and its standard deviation ($C_{max}$) of CP-Sat and our SPN512 on each shape. The last row (Gap sum from CP-Sat) offers a relative comparison by summing up the gaps of the SPN512 computed from the solutions produced by CP-Sat. A negative value means that the SPN produced better solutions.
As one can see, the SPN512 produces better average results than CP-Sat with just a fraction of time. On these 80 instances, we observed only 7 cases in which the solutions of CP-Sat are slightly better than those of the SPN. This remarks once more the quality of our proposal, making it a valid alternative to CP solvers for very large instances.
Lastly, we remark that on these very large instances, it is possible to further reduce the execution times of our SPN by employing speed-up techniques such as model quantization (i.e., converting floats to lower precision numbers) and the new compile functionality available in PyTorch $\geq 2.0$. Our model did not undergo any of the above (herein and in any other evaluation) and has been coded by keeping the implementation as simple as possible.
Appendix F Execution Times
Greedy Constructive Self-Labeling Pointer Network (SPN)
Shape PDRs INSA L2D CL $\beta$=1 $\beta$=32 $\beta$=64 $\beta$=128 $\beta$=256 $\beta$=512
$15\times 15$ 0.00 0.02 0.39 0.81 0.19 0.21 0.21 0.21 0.23 0.29
$20\times 15$ 0.00 0.03 0.60 1.10 0.28 0.28 0.28 0.28 0.36 0.49
$20\times 20$ 0.00 0.04 0.72 1.39 0.37 0.37 0.37 0.37 0.48 0.66
$30\times 15$ 0.01 0.08 0.95 1.49 0.40 0.41 0.41 0.47 0.61 0.92
$30\times 20$ 0.01 0.11 1.41 1.72 0.54 0.56 0.56 0.62 0.82 1.22
$40\times 15$ 0.01 0.14 1.45 1.81 0.55 0.56 0.56 0.68 0.93 1.45
$40\times 20$ 0.01 0.20 1.60 2.07 0.70 0.74 0.74 0.90 1.23 1.91
$50\times 15$ 0.01 0.23 1.81 2.82 0.80 0.79 0.83 0.92 1.32 2.15
$50\times 20$ 0.02 0.33 3.00 3.93 0.91 0.92 0.98 1.23 1.76 2.86
$100\times 20$ 0.19 1.61 9.39 9.58 2.10 2.16 2.49 3.54 5.81 9.35
Tot Sum 0.26 2.79 21.32 26.72 6.84 7.00 7.43 9.23 13.53 21.29
Figure 5: The average execution time in seconds of the coded algorithms on different instance shapes considered in Table 2. The times of PDRs, INSA, L2D, and CL refer to the construct of a single solution. Whereas we report the times of the SPN for varying numbers of sampled solution $\beta\in\{1,32,64,128,256,512\}$.
We additionally assess the timing factor, an important aspect in some scheduling scenarios. To this end, we plot in the left of Figure 5 the execution time trends of dispatching rules (PDRs), INSA, L2D (we omit CL as it has slightly higher execution times than L2D), and SPN on typical benchmark shapes, all applied in a greedy constructive manner. We also include the execution time of our SPN when sampling 512 solutions (SPN512) to prove that sampling multiple parallel solutions does not dramatically increase times. Note that all these algorithms were executed on the machine described at the beginning of Section 5. We omit other RL proposals as they do not publicly release the code, do not disclose execution times, or is unclear whether they used GPUs.
When sampling 512 solutions, the SPN512 takes less than a second on medium shapes, less than 3 sec on large ones, and around 10 sec on the big $100\times 20$. As these trends align with L2D, which only constructs a single solution, our SPN is a much faster neural alternative, despite being also better in terms of quality. We also underline that the SPN and any ML-based approach are likely to be slower than PDRs and constructive heuristics, either applied in a greedy or randomized way. One can see that such approaches work faster: PDRs take 0.2 sec and INSA 1.61 sec on $100\times 20$ instances. However, this increase in execution time is not dramatic and is largely justified by better performance, especially in the case of our SPN.
We finally provide in the right table of Figure 5 the precise SPN timing for varying numbers of sampled solutions $\beta$ during testing. Each row reports the average execution time on a benchmark shape and the last row (Tot Sum) sums up all the times for an overall comparison. We observe that sampling 32 and 64 solutions from the SPN takes the same except in large instances ($50\times 15,50\times 20$, and $100\times 20$ shapes), where variations are small. Therefore, sampling less than 32 solutions is pointless, as the reduction in the execution times is negligible while the quality is further reduced. Additionally, sampling more than 512 solutions further improves the SPN’s results at the cost of a consistent increment in the execution time. As an example, by sampling 1024 solutions, the overall average gap (Avg row in Table 2) is reduced by 0.4% in both benchmarks with respect to the $\beta$ = 512 case, but the execution time is roughly 1.9 times larger. Instead of blindly sampling many random solutions with long executions, it should be possible to achieve better results in less time with ad-hoc strategies, such as Beam-Search or other strategies, that sample differently based on the generated probabilities or other information. We leave this study on testing strategies to future work.
Appendix G Extended Results
The extended results of the coded algorithms (PDRs, INSA, SBH, L2D, MIP, CP, and the SPN’s configurations) are available at  [^2] Extended results: https://github.com/AndreaCorsini1/SelfLabelingJobShop/blob/main/output/Results.xlsx. For computing the percentage gap (PG), we used the optimal or best-known makespan of an instance ($C_{ub}$), available at: https://optimizizer.com/jobshop.php. The extended results highlight that models trained using our Self-Labeling strategy with $\beta<256$ exhibit competitive performance and, in some instances, even surpass the model trained with $\beta=256$. This indicates the efficacy of our strategy even with a reduced number of sampled solutions. However, as remarked in the paper, sampling a larger number of training solutions yields improved overall performance, as evidenced by the lower total sums observed for larger values of $\beta$.